OpenStudy (anonymous):

how do you find the axis of symmetry? heres and example equation: F(x)=3(x+4)^2+1

OpenStudy (jchick):

The axis of symmetry is the vertical line that goes through the vertex of a quadratic equation.

OpenStudy (anonymous):

@jchick can you help with this too? since its in vertex form would it be (4,1)?

OpenStudy (anonymous):

I understand how to get it from a graph just not in vertex or standard equation form

OpenStudy (jchick):

y=3(x+4)2+1

OpenStudy (jchick):

That is the vertex form

OpenStudy (jchick):

(−4,1) is the vertex

OpenStudy (anonymous):

oh okay, I hforget that the h spot hould be a -

OpenStudy (jchick):

The standard form of a parabola's equation is generally expressed: y=ax2+bx+c

OpenStudy (jchick):

The vertex form of a parabola's equation is generally expressed as: y = a(x-h)^2+k

OpenStudy (jchick):

If a is positive then the parabola opens upwards like a regular "U". If a is negative, then the graph opens downwards like an upside down "U".

OpenStudy (anonymous):

uh huh, and the axis is normally x=-b/2 a right ? for standard?>

OpenStudy (jchick):

Do you mean The axis of symmetry?

OpenStudy (anonymous):

yes

OpenStudy (jchick):

ok

OpenStudy (jchick):

Axis of Symmetry from Standard Form ax^2 + bx + c

OpenStudy (anonymous):

I get the vertex form now and what changes the direction a parabola open just not how to get the axis of symmetry in standard form.

OpenStudy (jchick):

a(x - h)^2 + k is vertex form

OpenStudy (jchick):

The first one is the standard form

OpenStudy (jchick):

Axis of Symmetry in Standard Form ax^2 + bx + c

OpenStudy (anonymous):

yeah. so what gets the axis from standard?

OpenStudy (jchick):

The standard form of a parabola's equation is generally expressed: y=ax2+bx+c The role of 'a' If a>0, the parabola opens upwards if a<0 it opens downwards.

OpenStudy (anonymous):

wait I thought that the standard form equation was ax^2+bx+C not the axis of symmetry .....

OpenStudy (jchick):